OpenStudy (anonymous):

HELP PLZZZ NEED IT NOW PLEASE HELPEvaluate the expression negative twelve minus six a minus four b for each a and b variable pair. Match the variable pair with the resulting value of the expression. the rest is below !!!

OpenStudy (anonymous):

a=1 and b=2 -14 2 -10 -26 a=-1 and b=-2 -14 2 -10 -26 a=-1 and b=2 -14 2 -10 -26 a=1 and b=-2 -14 2 -10 -26

OpenStudy (blurbendy):

I'm not sure exactly what you're asking